# From Internal Conflict to Contextual Adaptation of Language Models

### v1

@@ 6.5 Predictors of Persuasion
In our previous investigations, we do not see that the semantic uncertainty of a contextless question is a meaningful indicator of model persuasion. Furthermore, dataset partitions with anticipated increase in intra-memory conflict (i.e. temporal and disputable data) show lower persuasion than static facts. Therefore, we assess the strength of each relation to our CP score. We show the results of our linear regression test in Table 3. Though we do not find a very strong fit to the data, given its volatility ($R^{2}=0.185$), we do find several significant predictors. The strongest predictor is the number of edits of a fact, where it shows an inverse relationship to a fact’s persuasion score. This aligns with the behaviour we have seen in our comparison of temporal and static facts in Table 1.Furthermore, while object ($o_{pop}$) popularitysignificantly affects a fact’s persuasion score (previously shown by Du et al. (2024)) the magnitude of this effect is smaller than that of the number of edits of a fact or the semantic uncertainty of the contextless question ($SE_{q}$). Furthermore, the uncertainty of the contextless question alone is insufficient to predict a fact’s persuasion score, as determined by our correlation analysis in §6.3. In contrast, we see a significant correlation between the number of edits of a fact and the CP score ($r(2944)=-0.27,p=1.6e-42$). These findings typically hold across other models (see Appendix B), yet we find that the direction of the effect of $SE_{q}$ varies across models. Taken together, we can see that fact dynamicity, or intra-memory conflict, plays a bigger role in knowledge conflicts than fact popularity Mallen et al. (2023); Du et al. (2024). We also find an inverse effect of intra-memory conflict on the model’s susceptibility to persuasion for a given instance. Facts that change regularly are less likely to be updated with context-retrieval, yet facts that never change are easily persuaded.
Predictor $\widehat{\beta}$ $p$-value
Intercept 6.99(±$\pm 0.06$) $p<$2e-16*
# edits -0.65(±$\pm 0.06$) $p<$2e-16*
$o_{pop}$0.16(±$\pm 0.06$) 0.0106*
$s_{pop}$0.03(±$\pm 0.06$) 0.5743
$SE_{c}$1.16(±$\pm 0.06$) $p<$2e-16*
$SE_{q}$ 0.47(±$\pm 0.06$) 1.04e-13*
Table 3: The estimated coefficients ($\hat{\beta}$) and $p$-values of the linear model predicting the persuasion score. ($R^{2}=0.1848$).
@@ 7 Conclusion
We investigate the interaction between intra-memory conflict in context-memory conflicts, using multiple natural causes of intra-memory conflict (i.e. fact ‘dynamicity’), for the first time. While we approximate intra-memory conflict with number of fact edits and a semantic uncertainty score, we find that semantic uncertainty is a poor indicator of intra-memory conflict. For the context-memory conflicts, we propose the Coherent Persuasion score to measure the persuasiveness of the given context with consideration of the entire semantic output. Furthermore, we release DynamicQA, a QA dataset with realistic dynamic and static questions, (i.e. facts that naturally vary and those that do not) alongside additional context and realistic replacements for use in intra-memory and counter-memory conflict studies. We then evaluate three state-of-the-art LMs on their performance on static, temporal and disputable facts. We find that, surprisingly, static facts are the most easily updated with additional context, in comparison to temporal and disputable facts. Furthermore, our extensive analysis show that the number of unique presentations of the fact on Wikidata and Wikipedia (i.e. number of edits) has an inverse correlation with a model’s propensity to adapt that fact. We also find large model-level differences: Llama-2 is most easily persuaded, yet also the highest uncertainty. Over a variety of analyses, we find that uncertainty alone does not consistently indicate persuasion, requiring other approaches to update model knowledge than retrieval-augmentation in low-certainty domains.
@@
In our paper, we do not investigate the impact of model size on knowledge conflicts, as we were computationally limited to smaller models of a size of 7B parameters. We examine several different model architectures within the model size considered. We hypothesise that the effects observed might be less pronounced for larger models. We did not include models smaller than 7B, as the ones we tested showed poor performance on the QA task in initial experiments. Other 7B parameter models we investigated showed very low accuracy, even with provided context, which is why we chose to exclude them (Falcon-7B, Gemma-7B). Almazrouei et al. (2023); Team et al. (2024)
We have an imbalanced distribution of partitions in our dataset, owing to the difficulty of identifying disputable questions. We did our best to ensure our disputable dataset includes limited misinformation, though there is the possibility that some few instances made it through, given the difficulty of the task (also reflected in the relatively low annotator agreement). We also only use one possible measure of uncertainty. There have been a swarth of recent uncertainty and semantic consistency measures introduced for text generation Fadeeva et al. (2023), out of which we chose the most popular one.
@@ Acknowledgements
[Uncaptioned image][Uncaptioned $\end{array}$ This research was co-funded by the European Union (ERC, ExplainYourself, 101077481), by the Pioneer Centre for AI, DNRF grant number P1, as well as by The Villum Synergy Programme. We are also funded by the Villum and Velux Foundations Algorithms, Data and Democracy (ADD) grant, as well as the ERRATUM UCPH Data+ grant. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. We also thank Lovisa Hagström for her assistance in running experiments.
@@ A.1 Annotation
    Change. If both questions are incorrect or insufficient, the first question is edited. It must be rewritten so that most objects are acceptable answers to the question based on the context. If the question cannot be rewritten in a meaningful way, it should be discarded (i.e. if both words are adjectives). To ensure the text is specific enough, most extra information might be added. The annotator should also remove all "according to the text"s as this would not work well without context.
@@
    Discard. If there is no dispute for a question, it should be discarded. This could be if the two objects are two names for the same person (e.g. Kanye, Ye), two different spellings for the same word, or two synonyms (publications, writings). One exception is hypernyms (e.g. Danish, Scandinavian). We discard questions with insufficient context (e.g. ‘later jesus t.’). We remove fixed typos and vandalism and clear misinformation (e.g. the objects ‘china’ and ‘ukraine’ to the question ‘Which state agency is responsible for managing the Chernobyl exclusion zone and has offices on the site?’).
@@
Table 4: Example of prompt according to the presence of the context $c_{i}$ in the input. The example here is the question that asks about the producer of the movie Titanic with the answer-specific context which mentions answer (James Cameron) in the context.
@@
Table 5: Example of the prompt for the question generation. The LM is asked to generate the question about the KFC. And the expected answer of the question is “burger".
@@ Appendix B Additional Graphs
Here we present the results for additional models for the analyses in §6. We present the scatterplots for Llama-2 in Figure 6, and Qwen2 in Figure 7. Furthermore, we present the violin plots for Llama-2 in Figure 9, and Qwen in Figure 10. We present the results of the Linear Regression on Mistral persuasion score in Table 6 and Qwen2 persuasion score in Table 7.
Figure 5: Mistral instance level relationship between CP and SE.
Figure 6: Llama-2 instance level relationship between CP and SE.
Figure 7: Qwen2 instance level relationship between CP and SE.
Figure 8: Mistral SE distribution across partitions
Figure 9: Llama-2 SE distribution across partitions
Figure 10: Qwen2 SE distribution across partitions
Figure 11: Mistral CP distribution across partitions
Figure 12: Llama-2 CP distribution across partitions
Figure 13: Qwen2 CP distribution across partitions
Predictor $\widehat{\beta}$ $p$-value
Intercept 5.83(±$\pm 0.05$) $p<$2e-16*
# edits -0.37(±$\pm 0.05$) 4.71e-14*
$o_{pop}$0.27(±$\pm 0.05$) 1.68e-08*
$s_{pop}$0.15(±$\pm 0.05$) 0.00161*
$SE_{c}$0.06(±$\pm 0.05$) 0.17535
$SE_{q}$0.35(±$\pm 0.05$) 1.66e-12*
Table 6: Results of the Linear Regression model on Mistral Temporal Results. ($R^{2}=0.085$)
Predictor $\widehat{\beta}$ $p$-value
Intercept $\pm 0.04$) $p<$2e-16*
# edits -0.19(±$\pm 0.04$) 2.79e-06*
$o_{pop}$0.06(±$\pm 0.04$) 0.1223
$s_{pop}$0.10(±$\pm 0.04$) 0.0115
$SE_{c}$0.30(±$\pm 0.04$) 4.0e-15*
$SE_{q}$ 0.45(±$\pm 0.04$) $p<$2e-16*
Table 7: Results of the Linear Regression model on Qwen2 Temporal Results. ($R^{2}=0.083$)
Appendix C Additional Analyses
We further investigate the clustering behaviour we note in §6.3. We first show in Figure 14 that highly popular and temporal datapoints are not divided within the clusters.
Figure 14: We plot once again the instance-level relationship betwene persuasion and uncertainty with and without context, but now mark the location of the most temporal and popular facts.
We decide to linearly separate the data according to the seen behaviour by accepting instances with $SE_{c}<15.5$ as ‘class 0’, instances with $2116<SE_{c}<$ 21 as ‘class 1’, and instances with $22<SE_{c}$ as ‘class 2’. The classes are visualized in Figure 15. We show the 50 words with the top TF-IDF for each cluster questions and context in Table 8, where we find large overlaps in relevant words for the clusters. While Class 0 tends to have words more related to locations, and Class 1 includes words more related to bodies of work, we see words from both of these categories within Class 2, the most linearly separable cluster. Ultimately, this makes it difficult to deduce semantic causes in the question or context for the differential behaviour.
Figure 15: We show the identities of clusters 0, 1, and 2 on the persuasion vs. uncertainty scatterplot.
Class 0 questions Class 1 questions Class 2 questions
’festival’ ’france’ ’al’ ’turin’ ’jesus’ ’records’ ’assumption’ ’holiday’ ’grand’ ’let’ ’come’ ’general’ ’new’ ’lumpur’ ’david’ ’george’ ’invasion’ ’will’ ’charles’
’love’ ’city’ ’florence’ ’lima’ ’brisbane’ ’dublin’ ’london’  ’sport’ ’family’ ’world’ ’back’ ’fog’ ’day’ ’church’  ’le’ ’la’ ’space’ ’man’ ’michael’ ’nine’ ’born’ ’great’ ’you’ ’martin’
’japan’ ’grease’ ’taxi’ ’ho’ ’mohamed’  ’of’ ’the’ ’is’ ’friends’ ’here’ ’religion’ ’town’ ’life’ ’time’ ’dick’ ’sr’ ’kingdom’ ’play’ ’james’ ’sport’ ’brussels’
’orléans’ ’stockholm’ ’milan’ ’madrid’ ’italy’ ’composer’ ’country’ ’man’ ’and’ ’it’ ’in’ ’on’ ’love’ ’you’ ’author’ ’does’ ’england’ ’father’ ’richard’ ’country’ ’john’ ’robert’
’flag’ ’kingdom’ ’rome’ ’vienna’ ’religion’ ’jerusalem’ ’director’ ’capital’ ’composer’ ’genre’ ’what’ ’is’ ’author’ ’city’ ’religion’ ’william’ ’director’ ’composer’
’color’ ’occupation’   ’grape’ ’does’ ’celebration’ ’play’ ’occupation’ ’pilot’ ’to’ ’me’ ’end’ ’big’ ’my’  ’one’ ’club’ ’alex’ ’time’ ’what’ ’in’ ’the’ ’genre’ ’occupation’
’ ’in’ ’for’ ’screenwriter’ ’was’ ’who’ ’capital’ ’genre’ ’for’ ’screenwriter’ ’of’ ’was’ ’who’ ’halloween’ ’producer’ ’producer’ ’for’ ’capital’ ’screenwriter’ ’was’ ’who’ ’of’
Class 0 context Class 1 context Class 2 context
’second’ ’been’ ’first’ ’united’ ’world’ ’region’ ’after’ ’her’ ’best’ ’italian’ ’name’ ’french’ ’band’ ’had’ ’stars’ ’television’ ’after’ ’episode’
’episode’ ’he’ ’populous’ ’italian’ ’that’ ’one’ ’same’ ’first’ ’album’ ’that’ ’novel’ ’which’ ’co’ ’were’ ’their’ ’who’ ’produced’ ’based’ ’album’ ’
’became’ ’written’ ’million’ ’american’ ’metropolitan’ ’episode’ ’he’ ’also’ ’starring’ ’series’ ’who’ ’released’ ’released’ ’her’ ’its’ ’born’ ’that’ ’series’ ’also’
’directed’ ’at’ ’also’ ’film’ ’album’ ’most’ ’an’ ’which’ ’from’ ’based’ ’drama’ ’produced’ ’for’ ’an’ ’his’ ’which’ ’written’ ’from’ ’american’ ’for’
’centre’ ’known’ ’released’ ’has’ ’largest’ ’capital’ ’screenplay’ ’story’ ’comedy’ ’at’ ’its’ ’stars’ ’directed’ ’an’ ’known’ ’at’ ’as’ ’is’ ’by’
’from’ ’for’ ’population’ ’area’ ’its’ ’with’ ’on’ ’city’ ’to’ ’with’ ’written’ ’on’ ’as’ ’to’ ’was’ ’directed’ ’it’ ’his’ ’with’ ’on’ ’it’ ’he’ ’film’ ’to’ ’was’ ’one’
’was’ ’as’ ’it’ ’by’ ’is’ ’entity’ ’in’ ’and’ ’of’ ’in’ ’entity’ ’is’ ’of’ ’film’ ’by’ ’and’ ’american’ ’entity’ ’in’ ’of’ ’and’ ’she’ ’has’ ’city’
Table 8: Top 50 TF-IDF keywords for each context and question in the clusters identified in Figure 15.

### v2

@@ 6.4 Predictors of Persuasion
In our previous investigations, we see that the semantic entropy of a contextless question is not a meaningful indicator of model persuasion, though we do see stronger relationships for popularity and temporality. Furthermore, dataset partitions with anticipated increase in intra-memory conflict (i.e., temporal and disputable data) show lower persuasion than static facts. Therefore, we assess the strength of each relation to persuasion itself. We show the results of our logistic regression test in Table 3. Though we do not find a very strong fit to the data, given its volatility ($R^{2}=0.139$), we do find several significant predictors. The strongest predictor of persuasion is the number of edits of a fact, where it shows an inverse relationship to a fact’s persuasion. This aligns with the behaviour we have seen in our comparison of temporal and static facts in Table 1. Furthermore, while object popularity ($o_{pop}$) significantly affects a fact’s persuasion score (previously shown by Du et al. (2024)) the magnitude of this effect is smaller than that of the number of edits of a fact. These findings typically hold across other models (see Appendix B), where we show that $SE_{q}$ typically shows the weakest relation to model persuasion, and the number of edits to have one of the strongest relationships. Taken together, we can see that fact dynamicity, or intra-memory conflict, plays a bigger role in eliciting knowledge conflicts than fact popularity Mallen et al. (2023); Du et al. (2024). We also find an inverse effect of intra-memory conflict on an LM’s susceptibility to persuasion for a given instance. Facts that change regularly are less likely to be updated with context-retrieval, yet facts that never change are easily persuaded.
Predictor $\boldsymbol{\widehat{\beta}}$ $\mathbf{p}$-value
Intercept 0.87(±$\pm 0.008$) $p<$2e-16*
# edits -$\pm 0.008$) $p<$2e-16*
$o_{pop}$0.05(±$\pm 0.008$) 5.07e-9*
$s_{pop}$0.03(±$\pm 0.008$) 2.04e-5*
$SE_{c}$0.04(±$\pm 0.008$) 1.43e-6*
$SE_{q}$ 0.002(±$\pm 0.008$) 0.843
Table 3: The estimated coefficients ($\hat{\beta}$) and $p$-values of the linear model predicting the persuasion score. ($R^{2}=0.1386$). $\hat{\beta}$ values reflect the magnitude of each predictor’s effect on the dependent variable, the persuasion score, and the $p$-value denotes the statistical significance of the effect.
@@ 7 Conclusion
We investigate, for the first time, the effect of intra-memory conflict on context adapation, using multiple natural causes of intra-memory conflict (i.e., fact ‘dynamicity’). In this, we introduce a Coherent Persuasion score to measure the persuasiveness of the given context with consideration of the entire semantic output. Furthermore, we release DynamicQA, a QA dataset with realistic dynamic and static questions, (i.e., facts that naturally vary and those that do not) alongside additional context and realistic replacements for use in intra-memory and counter-memory conflict studies. We then evaluate three SOTA LMs on their performance on static, temporal and disputable facts. We find that, surprisingly, static facts are the most easily updated with additional context, in comparison to temporal and disputable facts. Furthermore, our extensive analyses show that the number of unique presentations of the fact on Wikidata and Wikipedia (i.e., number of edits) has an inverse correlation with a model’s propensity to adopt updates. We also find large model-level differences: Llama-2 is most easily persuaded. Over a variety of analyses, we find that uncertainty alone does not consistently indicate persuasion, requiring other approaches to update model knowledge than retrieval-augmentation in low-certainty domains.
@@
In our paper, we do not investigate the impact of model size on knowledge conflicts, as we were computationally limited to smaller models of a size of 7B parameters. We examine several different model architectures within the model size considered. We hypothesise that the effects observed might be less pronounced for larger models. We did not include models smaller than 7B, as the ones we tested showed poor performance on the QA task in initial experiments. Other 7B parameter models we investigated showed very low accuracy, even with provided context, which is why we chose to exclude them (Falcon-7B Almazrouei et al. (2023), Gemma-7B Team et al. (2024)).
We have an imbalanced distribution of partitions in our dataset, owing to the difficulty of identifying disputable questions. We did our best to ensure our disputable dataset includes limited misinformation, though there is the possibility that some few instances made it through, given the difficulty of the task (also reflected in the relatively low annotator agreement). We also only use one possible measure of uncertainty. There has been a swarth of recent uncertainty and semantic consistency measures introduced for text generation Fadeeva et al. (2023), out of which we chose the most popular one. We acknowledge relatively weak correlation values and $R^{2}$ reported in our paper; we owe these values to our use of realistic, rather than synthetic, retrieved contexts and distributions of popularity and temporality.
@@ Acknowledgements
[Uncaptioned image][Uncaptioned $\end{array}$ This research was co-funded by the European Union (ERC, ExplainYourself, 101077481), by the Pioneer Centre for AI, DNRF grant number P1, as well as by The Villum Synergy Programme. We are also funded by the Villum and Velux Foundations Algorithms, Data and Democracy (ADD) grant, as well as the ERRATUM UCPH Data+ grant. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. We also thank Lovisa Hagström for her assistance in running experiments and Nadav Borenstein and Pietro Tropeano for their careful proof-reading.
@@ A.1 Annotation
    Change. If both questions are incorrect or insufficient, the first question is edited. It must be rewritten so that most objects are acceptable answers to the question based on the context. If the question cannot be rewritten in a meaningful way, it should be discarded (i.e., if both words are adjectives). To ensure the text is specific enough, most extra information might be added. The annotator should also remove all "according to the text"s as this would not work well without context.
@@
    Discard. If there is no dispute for a question, it should be discarded. This could be if the two objects are two names for the same person (e.g., Kanye, Ye), two different spellings for the same word, or two synonyms (publications, writings). One exception is hypernyms (e.g., Danish, Scandinavian). We discard questions with insufficient context (e.g., ‘later jesus t.’). We remove fixed typos and vandalism and clear misinformation (e.g., the objects ‘china’ and ‘ukraine’ to the question ‘Which state agency is responsible for managing the Chernobyl exclusion zone and has offices on the site?’).
@@
Table 4: Example of a prompt according to the presence of the context $c_{i}$ in the input. The example here is the question that asks about the producer of the movie Titanic with the answer-specific context which mentions answer (James Cameron) in the context.
@@
Table 5: Example of a prompt for question generation. The LM is asked to generate the question about KFC. The expected answer of the question is “burger".
@@ Appendix B Additional Graphs
Here we present the results for additional models for the analyses in §6. Figure 4 shows that the analysis on the loss (§6.2) holds the same across the different models. Figures 5 and 6 show that temporality has a consistently negative and relatively strong relationship with the $CP$ score across models. While semantic entropy can also be a relatively strong correlate, it changes sign between models, showing that this relationship is inconsistent; meaning uncertain models may be more or less likely to utilise context. Tables 6 and 7 show that we continually see a strong effect of number of edits as a predictor of persuasion across models.
Predictor $\boldsymbol{\widehat{\beta}}$ $\mathbf{p}$-value
Intercept 0.88(±$\pm 0.01$) $p<$2e-16*
# edits -0.05(±$\pm 0.01$) 5.92e-12*
$o_{pop}$0.06(±$\pm 0.01$) 1.36e-13*
$s_{pop}$0.03(±$\pm 0.01$) 5.22e-4*
$SE_{c}$ 0.02(±$\pm 0.01$) 8.74e-3*
$SE_{q}$0.02(±$\pm 0.01$) 7.97e-3*
Table 6: Results of the Logistic Regression model on Mistral Temporal Results. ($R^{2}=0.116$)
Predictor $\boldsymbol{\widehat{\beta}}$ $\mathbf{p}$-value
Intercept 0.86(±$\pm 0.01$) $p<$2e-16*
# edits -0.07(±$\pm 0.01$) 4.60e-15*
$o_{pop}$0.06(±$\pm 0.01$) 1.64e-12*
$s_{pop}$0.04(±$\pm 0.01$) 1.21e-4*
$SE_{c}$ 0.08(±$\pm 0.01$) $p<$2e-16*
$SE_{q}$ 0.03(±$\pm 0.01$) 1.40e-4*
Table 7: Results of the Logistic Regression model on Qwen2 Temporal Results. ($R^{2}=0.197$)
Figure 4: The distribution of losses for each partition of the dataset on Mistral and Qwen
Figure 5: The instance-level relationship between Coherent Persuasion Score and 3 possible factors that impact persuasion: semantic entropy, subject popularity and object temporality, alongside their Pearson correlation scores. Temporality shows the strongest relationship with persuasion. We also highlight two behaviours of interest: persuaded and stubborn instances. These results are for Mistral.
Figure 6: The instance-level relationship between Coherent Persuasion Score and 3 possible factors that impact persuasion: semantic entropy, subject popularity and object temporality, alongside their Pearson correlation scores. Temporality shows the strongest relationship with persuasion. We also highlight two behaviours of interest: persuaded and stubborn instances. These results are for Qwen.